OpenStudy (anonymous):

will give medal!(:

OpenStudy (anonymous):

@jim_thompson5910 @ingenuus

OpenStudy (anonymous):

This is easy. Trust me. You don't need us...

OpenStudy (anonymous):

Basically its the diameter divided by 2.

jimthompson5910 (jim_thompson5910):

You already asked this same question? Am I missing something?

OpenStudy (anonymous):

no here! Sorry I meant to post another one @jim_thompson5910 1 sec

jimthompson5910 (jim_thompson5910):

sry, the site is being slow, one sec

jimthompson5910 (jim_thompson5910):

ok i see it, were you able to get anywhere with this one?

OpenStudy (anonymous):

well I think that the coordinates are x(-6) and y(7) ? (:

jimthompson5910 (jim_thompson5910):

good, the other two coordinates are what?

OpenStudy (anonymous):

that's what I'm stuck on :( Are they simular?

jimthompson5910 (jim_thompson5910):

the origin is the point (?, ?)

OpenStudy (anonymous):

i did this @jim_thompson5910 @ryan123345 a^2+b^2=c^2 6^2=7^2=c^2 36+49=c^2 sqrt. of 85=c=9.21 (:

OpenStudy (ryan123345):

9.22

OpenStudy (ryan123345):

btw

OpenStudy (ryan123345):